Circular No. 5918

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505

MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET KUSHIDA (1994a)

S. Nakano, Sumoto, Japan, reports that Yoshio Kushida has discovered a new comet on Technical Pan 6415 film. The following observations have been reported:

1994 UT  α_{2000}  δ_{2000}  m_{1} Observer
Jan. 8.80868 9^{h}45^{m}52.^{s}0 +3^{o}10′43′′ 13.5 Kushida
8.84896 9 45 52.8 +3 10 17  _{”}
9.52853 9 46 19.22 +3 03 23.8 12.5 Kojima
9.54000 9 46 19.59 +3 03 16.8  _{”}
9.54728 9 46 19.82 +3 03 12.3  _{”}
9.55382 9 46 20.24 +3 03 08.4 12.5 Urata
9.56111 9 46 20.27 +3 03 03.6  _{”}
9.56366 9 46 20.44 +3 03 02.6 11.9 Nakamura
9.56875 9 46 20.59 +3 02 59.5  _{”}
9.57431 9 46 20.76 +3 02 56.0  _{”}
9.57663 9 46 20.75 +3 02 53.3  Kobayashi
9.58205 9 46 20.90 +3 02 49.8  _{”}
9.58569 9 46 21.00 +3 02 47.7  _{”}
9.64286 9 46 22.73 +3 02 13.8 13.2 Kawanishi
9.64601 9 46 22.90 +3 02 12.1  _{”}
9.64899 9 46 22.93 +3 02 09.3  _{”}

Y. Kushida (Yatsugatake South Base Observatory). 0.10-m f/4.0 patrol camera. Comet has strong central condensation with coma diameter ∼ 1′- 2′; object near edge of patrol film, making positions somewhat uncertain.

T. Kojima (YGCO Chiyoda Observatory). 0.25-m f/6 reflector + CCD. Comet is diffuse with strong condensation, diameter 35′′; short tail extends to northwest.

T. Urata (Nihondaira Observatory). 0.31-m f/4.7 reflector + CCD (3- to 5-min exposures). Comet has tail 6′-10′ long toward northwest; coma diameter ∼ 60′′.

A. Nakamura (Kuma Kogen). 0.60-m f/6.0 Ritchey-Chrétien + CCD. Coma

diameter ∼ 2′.6. T. Kobayashi (Oizumi). 0.25-m f/4.4 reflector + CCD. K. Kawanishi (Akou). 0.33-m f/3.3 reflector + CCD.

1994 January 9 Daniel W. E. Green